# TimeSeries2Report prompting enables adaptive large language model management of lithium-ion batteries

### v1

@@ Abstract
Large language models (LLMs) offer promising capabilities for interpreting multivariate time-series data, yet their application to real-world battery energy storage system (BESS) operation and maintenance remains largely unexplored. Here, we present TimeSeries2Report (TS2R), a prompting framework that converts raw lithium-ion battery operational time-series into structured, semantically enriched reports, enabling LLMs to reason, predict, and make decisions in BESS management scenarios. TS2R encodes short-term temporal dynamics into natural language through a combination of segmentation, semantic abstraction, and rule-based interpretation, effectively bridging low-level sensor signals with high-level contextual insights. We benchmark TS2R across both lab-scale and real-world datasets, evaluating report quality and downstream task performance in anomaly detection, state-of-charge prediction, and charging/discharging management. Compared with vision-, embedding-, and text-based prompting baselines, report-based prompting via TS2R consistently improves LLM performance in terms of across accuracy, robustness, and explainability metrics. Notably, TS2R-integrated LLMs achieve expert-level decision quality and predictive consistency without retraining or architecture modification, establishing a practical path for adaptive, LLM-driven battery intelligence.
@@ 1 Introduction
From portable electronics [liang2019review] to electric vehicles [cano2018batteries, harper2019recycling, liu2025multi] and grid-scale renewable energy storage systems [arteaga2017overview, jiang2025battery], lithium-ion batteries (LIBs) power nearly every facet of modern life. As deployment scales from individual cells to gigawatt-hour battery energy storage systems (BESS), ensuring the safety, reliability, and efficiency of such large, heterogeneous networks has become increasingly complex and mission-critical. Effective battery management requires precise, real-time monitoring of cell states including state of charge (SOC) [adaikkappan2022modeling, shen2017co], state of health (SOH) [hu2020battery, jones2022impedance, severson2019data], and temperature [cao2025model, mei2023operando] under highly dynamic load conditions, evolving degradation pathways, and environmental variability. Yet the coupled electrochemical, thermal, and mechanical dynamics across thousands of interacting cells make predictive maintenance and adaptive control fundamentally challenging [mama2025comprehensive, che2022state, xie2022faults].
@@
Large language models (LLMs) [naveed2025comprehensive, li2025self, dagdelen2024structured, guo2025deepseek, achiam2023gpt, zhao2025align, kevian2024capabilities] offer a compelling alternative, enabling flexible reasoning, contextual understanding, and natural language generation across diverse tasks. Their general-purpose nature opens the door to a new paradigm in battery management – one centered on explainability, adaptability, and task generalization. However, mainstream LLMs such as ChatGPT [achiam2023gpt] are designed primarily for text-based conversations and lack native support for interpreting continuous, multivariate, and temporally structured time-series data. As highlighted in recent studies [jiang2024empowering, wang2025itformer, 11268252], this architectural mismatch limits their utility in time-series-dominant domains such as battery OM, where key tasks like SOH estimation and fault diagnosis depend on rich temporal dynamics. Recent developments in multimodal LLMs, such as ChatTS [xie2024chatts] and Time-LLM [jin2023time], address this gap by learning shared embedding spaces for time-series and text, allowing LLMs to treat signal segments as pseudo-tokens. However, these models require extensive retraining and deep natural language processing (NLP) expertise, which limits their accessibility and practical use in the battery domain.
@@ 2.1 Illustration of the TimeSeries2Report framework
The generated report serves as readily usable inputs for downstream LLM-driven OM tasks. We demonstrate their applicability in three representative scenarios: SOC prediction, operational monitoring (abnormality detection), and charging/discharging management (Fig. 1d and Methods 4.8). Given these sentences and the task-specific user queries, the LLM generates interpretable responses supported by explicit, semantically grounded reasoning, thereby enhancing transparency and explainability throughout the battery OM workflow.
@@ Acknowledgements
Conceptualisation, J.Y. and C.Z.; Methodology, J.Y. and C.Z.; Software, J.Y. and C.Z.; Original data acquisition, J.Y. and C.Z.; Analysis, J.Y., C.Z., M.G. and Z.C.; Visualisation, J.Y. and C.Z.; Writing original draft, J.Y, C.Z., M.G. and Z.C.; Writing, review & editing, J.Y, C.Z., M.G. and Z.C.; Supervision, C.Z. and Z.C.

### v2

@@ Abstract
Large language models (LLMs) offer promising capabilities for interpreting multivariate time-series data, yet their application to real-world battery energy storage system (BESS) operation and maintenance remains largely unexplored. Here, we present TimeSeries2Report (TS2R), a semantic translation framework that converts raw lithium-ion battery operational time-series into structured, semantically enriched reports, enabling LLMs to reason, predict, and make decisions in BESS management scenarios. TS2R encodes short-term temporal dynamics into natural language through a combination of segmentation, semantic abstraction, and rule-based interpretation, effectively bridging low-level sensor signals with high-level contextual insights. We benchmark TS2R across both lab-scale and real-world datasets, evaluating report quality and downstream task performance in anomaly detection, state-of-charge prediction, and charging/discharging management. Compared with vision-, embedding-, and text-based prompting baselines, report-based prompting via TS2R consistently improves LLM performance in terms of across accuracy, robustness, and explainability metrics. Notably, TS2R-integrated LLMs achieve expert-level decision quality and predictive consistency without retraining or architecture modification, establishing a practical path for adaptive, LLM-driven battery intelligence.
@@ 1 Introduction
From portable electronics [liang2019review] to electric vehicles [chen2019recycling, harper2019recycling, li2019data] and grid-scale renewable energy storage systems [arteaga2017overview, jiang2025battery], lithium-ion batteries (LIBs) power nearly every facet of modern life. As deployment scales from individual cells to gigawatt-hour battery energy storage systems (BESS), ensuring the safety, reliability, and efficiency of such large, heterogeneous networks has become increasingly complex and mission-critical. Effective battery management requires precise, real-time monitoring of cell states including state of charge (SOC) [adaikkappan2022modeling, shen2017co], state of health (SOH) [hu2020battery, rhyu2025systematic, severson2019data], and temperature [cao2025model, mei2023operando] under highly dynamic load conditions, evolving degradation pathways, and environmental variability. Yet the coupled electrochemical, thermal, and mechanical dynamics across thousands of interacting cells make predictive maintenance and adaptive control fundamentally challenging [mama2025comprehensive, che2022state, xie2022faults].
@@
Large language models (LLMs) [naveed2025comprehensive, li2025self, zuo2025large, majumder2024exploring, achiam2023gpt, zhao2025align] offer a compelling alternative, enabling flexible reasoning, contextual understanding, and natural language generation across diverse tasks. Their general-purpose nature opens the door to a new paradigm in battery management – one centered on explainability, adaptability, and task generalization. However, mainstream LLMs such as ChatGPT [achiam2023gpt] are designed primarily for text-based conversations and lack native support for interpreting continuous, multivariate, and temporally structured time-series data. As highlighted in recent studies [jiang2024empowering, wang2025itformer, 11268252], this architectural mismatch limits their utility in time-series-dominant domains such as battery OM, where key tasks like SOH estimation and fault diagnosis depend on rich temporal dynamics. Recent developments in multimodal LLMs, such as ChatTS [xie2024chatts] and Time-LLM [jin2023time], address this gap by learning shared embedding spaces for time-series and text, allowing LLMs to treat signal segments as pseudo-tokens. However, these models require extensive retraining and deep natural language processing (NLP) expertise, which limits their accessibility and practical use in the battery domain.
@@ 2.1 Illustration of the TimeSeries2Report framework
The generated report serves as readily usable inputs for downstream LLM-driven OM tasks. We demonstrate their applicability in three representative scenarios: SOC prediction, operational monitoring (abnormality detection), and charging/discharging management (Fig. 1d and Methods 4.8). Given these sentences and the task-specific user queries, the LLM generates interpretable responses supported by explicit, semantically grounded reasoning, thereby enhancing transparency and explainability throughout the battery OM workflow. See supplementary video demo.mp4 for full demonstration.
@@ Acknowledgements
Conceptualisation, J.Yang and C.Zhao; Methodology, J.Yang and C.Zhao; Software, J.Yang and C.Zhao; Original data acquisition, J.Yang and C.Zhao; Analysis, J.Yang, C.Zhao, M.Guay and Z.Cao; Visualisation, J.Yang and C.Zhao; Writing original draft, J.Yang, C.Zhao, M.Guay and Z.Cao; Writing, review & editing, J.Yang, C.Zhao, M.Guay and Z.Cao; Supervision, C.Zhao and Z.Cao.